OpenStudy (anonymous):

Which of the following equations is parallel to the line passing through A(-2,7), B(-2,-2)

OpenStudy (anonymous):

well first let's find the equation with those points y=mx+b m=(y2-y2)/(x2-x1) m=(7+2)/(-2+2) m=9/0 m=undefined if m is undefined it would be x=# so the equation is x=-2